### v1

@@ 1 Introduction and Motivation
To bridge this critical gap, this paper introduces Sim4IA-Bench, the first public benchmark resource specifically designed for the evaluation of user simulators. As the primary contribution of this work, we introduce a dataset derived from a recent user simulation initiative, the Sim4IA [36] Micro-Shared Task tackling interactive IR simulation (Task A) and conversational session simulation (Task B). In addition to submissions from participating approaches, we complement this dataset with a proposed set of string-based and system-based similarity measures, offering a crucial starting point for the community to assess simulator quality. Importantly, our goal is not to measure a simulator’s success in a downstream retrieval task, but to directly address the more fundamental question of how well it reproduces authentic user behavior.
Beyond the methodological gap, a significant practical barrier has also hinders the wider adoption of user simulation: the substantial infrastructure and engineering effort required to build a simulator from scratch. To address this challenge, Sim4IA-Bench provides a comprehensive suite of practical resources designed to dramatically lower this barrier to entry. At the core, it includes a simulation toolkit that serves as a starting kit with baseline implementations, data loaders, and evaluation scripts. Sim4IA-Bench contains a rich collection of artifacts, such as prepared session logs, participants’ run files, and comprehensive documentation detailing data formats and evaluation protocols. By packaging these components together, we shift the focus from foundational engineering and enable researchers to concentrate on the core scientific challenges of simulator design. Sim4IA-Bench [^1] GitHub repository of Sim4IA-Bench: https://github.com/irgroup/Sim4IA-Bench is released under the MIT license, enabling both academic and industry researchers to access and use the resource.
@@ 1.0.1 The Sim4IA-Bench Suite
Furthermore, this work is intended to guide the development of future community-wide evaluation initiatives. The methodology, dataset structure, and experiences gained from organizing this shared task provide a valuable blueprint for establishing larger-scale, standardized evaluation campaigns at TREC or CLEF, and Sim4IA-Bench will be maintained as part of the User Simulation subtask (Task 3) in LongEval@CLEF’26.
@@ 2.1 Validating User Simulators
While simulators need to be validated against real human interactions [3], the specific requirements for simulators differ depending on what they are going to be used for, i.e. training vs. evaluation [4]. In general, such a comparison against human interactions may be performed at a distributional level, for example, by comparing (i) query characteristics (length, terms) [2] or similarity [21] for traditional search or (ii) the distribution of dialogue acts or success rate for conversational agents [52]. Other approaches for validation include labeling specific instances according to different dimensions, like naturalness, usefulness, grammar for conversational utterances [46, 37, 54] or comparing entire conversations (human vs. simulated) in a side-by-side manner [52, 46]. Another method is the evaluation based on testers where testers are sets of IR systems over which a specific performance pattern can be expected that a simulator is trying to reproduce [23, 24].
While there are few resources dedicated to validate simulators, with Sim4IA-Bench we provide exactly this to bridge this gap.
Table 1: Comparison of Interactive IR and Conversational Search datasets for next-query or next-utterance prediction. Size refers to the number of sessions or conversations in each dataset. T informs whether the dataset covers traditional IR, C informs whether the dataset covers conversations. A refers to additional public assets like system runs from a shared task. Domain indicates the topic of the dataset.
Dataset Size T C A Domain AOL [33] 283,207 (AOL17) Web search SUSS [28] 484,449 Academic search Yandex [39] 797,867 Web search TREC Session [8] 1564 Web search TianGong-ST [10] 147,155 Web search TripClick [34] 1,602,648 Health Baidu-ULTR [57] 1.2 bil Web search Webis-FUQ-24 [21] 18,980 Web search, arguments, exhibitions, product search Persona-Chat [53] 10,907 Human chatting Webis-CQR-2 [20] 284 Arguments, books, news, trips SoguoQ [40] 14,075,717 Web search TREC CAsT (2022) [31] 50 Web search LLM-REDIAL [25] 47,600 Movies, books, sports WildChat [55] 1,039,785 Web search LMSYS-CHAT-1M [56] 1,000,000 Web search Ours 160 Academic search
@@ 2.2 Datasets
There are several large-scale datasets that primarily originate from the web search domain or rather small-scale domain-specific datasets from academic, health-related, and other fields. Most of these datasets serve as evaluation toolkits for different aspects of the user modeling in an interactive search setting with varying degrees of an explicit user simulation. For instance, some datasets allow a comprehensive evaluation of the different interactions in a simulated sessions, while other have a more specific focus like query suggestion, click modeling, or utterance prediction. Building on these datasets, prior work has examined related aspects such as query expansion and suggestion [29, 12], query and utterance prediction [44, 18, 27, 48], or session modeling [13, 9, 49, 17].
@@
Taken together, these datasets exemplify how complementary assets facilitate structured comparison, either by providing directly comparable shared-task runs or by being integrated into established benchmarks. Building on this, our resource is the first to provide complementary assets specifically for next query prediction in the context of user simulation, offering run files that enable reproducible system comparisons and systematic evaluation of measures under controlled experimental conditions
@@ 3.1 Task Descriptions and Dataset Contents
We provide session-based datasets for the Sim4IA Micro-Shared Task [36], designed to support the evaluation and development of user simulators in interactive IR. The resources include two sets of data corresponding to the two task variants: Task A (interactive IR simulation) and Task B (conversational session simulation). Each task has a training set of 45 sessions and a test set of 35 sessions. The test sets differ from the training sets in that the final query of each session is withheld, allowing participants to evaluate simulator predictions without overfitting. In all tasks ten next queries or utterances are to be predicted.
For a concise overview of the two tasks and their respective workflows, see Figure 1.
@@ 3.1.1 Task A: Interactive IR Simulation
Each session in Task A contains queries, the corresponding retrieved SERPs, and the documents clicked by users. Clicks are categorized into three types: clicks on authors, clicks on the work itself, and clicks on the “Download PDF” option. Timestamps for queries and clicks are included, enabling simulators to account for temporal aspects of interactions and to model different types of user behavior. We composed two variants of Task A using the same data: Task A1 only considers the last query and corresponding SERP information while in Task A2 the whole session was allowed to be used.
@@ 3.1.2 Task B: Conversational Session Simulation
Task B sessions contain only utterance-response pairs. Responses were generated using Google’s Gemma 3 12B model served on Ollama. For each query, the model was provided with the top three retrieved documents, the preceding response, and all upcoming utterances. The prompt (see Figure 2) was designed to ensure that the model focused solely on the current utterance while steering towards the upcoming utterances, producing coherent and plausible conversational flows.
Utterance: What are the main symptoms of diabetes? Relevant Documents: Title: Diabetes Overview
@@
Previous Response: Diabetes is a condition affecting the body’s ability to regulate blood sugar. Upcoming User Utterances: Utterance: How can diabetes be managed effectively?
@@
Instruction: - Answer the query using the relevant documents and the previous response. - Act like a RAG system: provide relevant, informative, and context-aware responses tailored to the user’s query intent. - Try to only answer the current utterance and explicitly do provide answers in the response, that might answer the provided upcoming utterances. - Limit your answer to no more than 150 words. - Focus on key points and avoid unnecessary repetition. - Please do answer formally and don’t use phrases like ’Okay, here’s a response to “Diabetes”’ acting as a RAG system. Answer:
Figure 2: Prompt template for generating the synthetical responses of the conversational system for Task B. Black parts show the unchanging structural and instructional parts, purple components depict the variables inserted depending on sessions and utterances.
3.1.3 Dataset Statistics
For Task A the dataset contains an average session length of 5.20 queries (4.20 reformulations on average), with 1.49 clicks per query. For Task B the average session length is 4.85 queries (3.85 reformulations on average).
@@ 3.2 Session Extraction from CORE Logs
Sessions were reconstructed from CORE [22] log files using a heuristic approach to group queries into sessions: starting from the last query in a potential session, queries occurring within a time window of -10 to +5 minutes were considered candidates for inclusion. Queries were added to a session if their cosine similarity, calculated using embeddings generated by the SentenceTransformer model all-MiniLM-L6-v2, with the current session queries was greater than or equal to 0.1. All reconstructed sessions were manually reviewed to ensure coherence and plausibility. Query sessions that contained fewer than three reformulations or did not originate from a plausible information need were excluded.
@@ 3.4 Submitted Runs
For the Micro-Shared Task we obtained 62 runs over all tasks in total and three accompanying lab notes detailing information on these runs by groups from CIR [7], Webis [16] and THM [15]. In general we distinguish between runs which were composed (semi-) manually, through an LLM as well as rule-based ones. We further differentiate LLM-based runs as ones composed by prompting the LLM to behave as a persona alone, through a combination of tuning and prompting and others. Table 2 provides the number of runs of different types from each group for the respective tasks.
@@ 4 Assessing Simulator Fidelity
Evaluating the quality and validity of a simulator in query prediction tasks is challenging, and the choice of suitable measures remains an open question. In this work, our methodology shifts the focus from a simulator’s effectiveness in retrieval tasks to its reproduction quality—that is, how well it replicates the authentic user behavior observed in real interaction logs. To assess this, we employ a complementary set of string-based and system-based similarity measures. Each measure is designed to capture a different aspect of simulator fidelity by quantifying the degree to which simulated queries match actual user inputs.
@@ 4.1 Measures
Semantic Similarity. Semantic similarity assesses how close the predicted queries are in meaning to the original user queries. We computed the average over $N$ sessions of the cosine similarity between sentence embeddings of the original query $q_{i}^{true}$ and the $K$ candidate queries $q_{i,j}$, bounding the similarity values from [1,1][-1,1] to $[0,1]$ for better visual comparability with the other measures.
$\bar{S}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{K}\sum_{j=1}^{K}\text{cosine}(q_{i}^{true},q_{i,j})\right)$
@@
Redundancy. To evaluate novelty and diversity, we introduced a “Redundancy” measure, which calculates the Jaccard similarity between all $K$ candidate queries or utterances within a session averaged across all $N$ sessions:
$\bar{R}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{2}{K(K-1)}\sum_{1\leq j<k\leq K}\text{Jaccard}(q_{i,j},q_{i,k})\right)$
Low redundancy indicates that a simulator produces multiple distinct candidates that remain semantically similar to the original query, whereas high redundancy signals minimal variation.
SERP Overlap. To capture system-level effects, we measured the average overlap between the search engine results retrieved for the original query and the $K$ candidates over the $N$ sessions:
$\bar{O}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{K}\sum_{j=1}^{K}\text{overlap}(q_{i}^{\text{true}},q_{i,j})\right)$
This measure provides insight into how the predicted queries would influence retrieval outcomes, complementing the string- and meaning-based evaluations.
@@
$\text{RDS}_{\cos\geq 0.7}=\frac{1}{|S|}\sum_{s\in S}\left(\sum_{q\in Q_{s}}\frac{1}{\text{rank}_{q}^{\cos\geq 0.7}}\right)$
RDS rewards simulators that rank high-quality, diverse candidate queries or utterances at the top while penalizing poor ordering and runs submitting fewer than the required ten candidates, as this lowers their potential multiplier.
@@ 4.2 Case Study: Analysis of Task A1
(a) Semantic Similarity
(b) SERP Overlap
(c) Redundancy
(d) Rank-Diversity Score
@@ 5 Reflections on the Shared Task
The considerable effort invested in preparing the study environment, including comprehensive setup instructions and video guides, proved successful. None of the participating teams required substantial organizer support, demonstrating the effectiveness and reusability of the provided resources. The smooth onboarding process was a key logistical success, ensuring that the setup can be easily reproduced in future studies.
The shared task and its resulting dataset provide a valuable first step toward systematically exploring user simulation in interactive search. However the process also highlighted several key challenges and limitations. A significant finding, particularly for Task B (conversational session simulation), was that the simulated utterances remained largely “query-like” and lacked the natural verbosity typical of this setting. This points to a major area for future improvement in simulator design. Furthermore, the semi-synthetic nature of the dataset highlighted both possibilities and constraints, offering opportunities for experimentation while revealing areas where realism could be improved. Finally, the task underscored that selecting appropriate evaluation measures is still an open and underexplored challenge for the community.

### v2

@@ 1 Introduction and Motivation
To bridge this critical gap, this paper introduces Sim4IA-Bench, the first public benchmark resource specifically designed to evaluate user simulators. As the primary contribution of this work, we introduce a dataset derived from a recent user simulation initiative, the Sim4IA [43] Micro-Shared Task, tackling interactive IR simulation (Task A) and conversational session simulation (Task B). In addition to submissions from participating approaches, we complement this dataset with a proposed set of string-based and system-based similarity measures, offering a crucial starting point for the community to assess simulator quality. Importantly, our goal is not to measure a simulator’s success in a downstream retrieval task, but to directly address the more fundamental question of how well it reproduces authentic user behavior.
Beyond the methodological gap, a significant practical barrier has also hindered the wider adoption of user simulation: the substantial infrastructure and engineering effort required to build a simulator from scratch. To address this challenge, Sim4IA-Bench provides a comprehensive suite of practical resources designed to dramatically lower this barrier to entry. At the core, it includes a simulation toolkit that serves as a starting kit with baseline implementations, data loaders, and evaluation scripts. Sim4IA-Bench contains a rich collection of artifacts, such as prepared session logs, participants’ run files, and comprehensive documentation detailing data formats and evaluation protocols. By packaging these components together, we shift the focus from foundational engineering and enable researchers to concentrate on the core scientific challenges of simulator design. Sim4IA-Bench [^1] GitHub repository of Sim4IA-Bench: https://github.com/irgroup/Sim4IA-Bench is released under the MIT license, enabling both academic and industry researchers to access and use the resource.
@@ 1.0.1 The Sim4IA-Bench Suite
Furthermore, this work is intended to guide the development of future community-wide evaluation initiatives. The methodology, dataset structure, and experiences gained from organizing this shared task provide a valuable blueprint for establishing larger-scale, standardized evaluation campaigns at TREC or CLEF, and Sim4IA-Bench will be maintained as part of the User Simulation subtask (Task 3) in LongEval@CLEF’26 [6].
@@ 2.1 Validating User Simulators
While simulators need to be validated against real human interactions [4], the specific requirements for simulators differ depending on what they are going to be used for, i.e. training vs. evaluation [5]. In general, such a comparison against human interactions may be performed at a distributional level, for example, by comparing (i) query characteristics (length, terms) [3], similarity [26], or retrieval performance and shared task utility [7] for traditional search or (ii) the distribution of dialogue acts or success rate for conversational agents [60]. Other approaches for validation include labeling specific instances according to different dimensions, like naturalness, usefulness, grammar for conversational utterances [53, 44, 61] or comparing entire conversations (human vs. simulated) in a side-by-side manner [60, 53]. Another method is the evaluation based on testers where testers are sets of IR systems over which a specific performance pattern can be expected that a simulator is trying to reproduce [29, 30].
While there are few resources dedicated to validating simulators, with Sim4IA-Bench we provide exactly this to bridge this gap.
Table 1: Comparison of Interactive IR and Conversational Search datasets for next-query or next-utterance prediction. Size refers to the number of sessions or conversations. T informs whether the dataset covers traditional IR, C informs whether the dataset covers conversations. A refers to additional public assets like system runs from a shared task. Domain indicates the topic of the dataset.
Dataset Size T C A Domain AOL [40] 283,207 (AOL17) Web search SUSS [35] 484,449 Academic search Yandex [46] 797,867 Web search TREC Session [12] 1564 Web search ConvAI [34] 4750 Human Chatting TianGong-ST [14] 147,155 Web search ConvAI2 [20] 4406 Human Chatting TripClick [41] 1,602,648 Health ConvAI3 (ClariQ) [1] 1,596,757 Human Chatting Baidu-ULTR [64] 1.2 bil Web search Webis-FUQ-24 [26] 18,980 Web search, arguments, exhibitions, product search Persona-Chat [59] 10,907 Human chatting Webis-CQR-2 [25] 284 Arguments, books, news, trips SoguoQ [47] 14,075,717 Web search TREC CAsT (2022) [38] 50 Web search LLM-REDIAL [31] 47,600 Movies, books, sports WildChat [62] 1,039,785 Web search LMSYS-CHAT-1M [63] 1,000,000 Web search Ours 160 Academic search
@@ 2.2 Datasets
There are several large-scale datasets that primarily originate from the web search domain or rather small-scale domain-specific datasets from academic, health-related, and other fields. Most of these datasets serve as evaluation toolkits for different aspects of the user modeling in an interactive search setting with varying degrees of an explicit user simulation. For instance, some datasets allow a comprehensive evaluation of the different interactions in a simulated sessions, while others have a more specific focus like query suggestion, click modeling, or utterance prediction. Building on these datasets, prior work has examined related aspects such as query expansion and suggestion [36, 16], query and utterance prediction [51, 23, 33, 55], or session modeling [17, 13, 56, 22].
@@
Taken together, these datasets exemplify how complementary assets facilitate structured comparison, either by providing directly comparable shared task runs or by being integrated into established benchmarks. Building on this, our resource is the first to provide complementary assets specifically for next query prediction in the context of user simulation, offering run files that enable reproducible system comparisons and systematic evaluation of measures under controlled experimental conditions.
@@ 3.1 Task Descriptions and Dataset Contents
We provide session-based datasets for the Sim4IA Micro-Shared Task [43], designed to support the evaluation and development of user simulators in interactive IR. The resources include two sets of data corresponding to the two task variants: Task A (interactive IR simulation) and Task B (conversational session simulation). Each task has a training set of 45 sessions and a test set of 35 sessions. The test sets differ from the training sets in that the final query of each session is withheld, allowing participants to evaluate simulator predictions without overfitting. In all tasks ten next queries or utterances are to be predicted. For a concise overview of the two tasks and their respective workflows, see Figure 1.
@@ 3.1.1 Task A: Interactive IR Simulation
Each session in Task A contains queries, the corresponding retrieved SERPs, and the documents clicked by users. Clicks are categorized into three types: clicks on authors, clicks on the work itself, and clicks on the “Download PDF” option. Timestamps for queries and clicks are included, enabling simulators to account for temporal aspects of interactions and to model different types of user behavior. We composed two variants of Task A using the same data: Task A1 only considers the last query and corresponding SERP information, while in Task A2, the whole session was allowed to be used. For Task A, the dataset contains an average session length of 5.20 queries (4.20 reformulations on average), with 1.49 clicks per query.
@@ 3.1.2 Task B: Conversational Session Simulation
Task B sessions contain only utterance-response pairs. Responses were generated using Google’s Gemma 3 12B model served on Ollama. For each query, the model was provided with the top three retrieved documents, the preceding response, and all upcoming utterances. The prompt (see Figure 2) was designed to ensure that the model focused solely on the current utterance while steering towards the upcoming utterances, producing coherent and plausible conversational flows. For Task B, the average session length is 4.85 queries (3.85 reformulations on average).
Utterance: What are the main symptoms of diabetes?
Relevant Documents:
Title: Diabetes Overview
@@
Previous Response: Diabetes is a condition affecting the body’s ability to regulate blood sugar.
Upcoming User Utterances:
Utterance: How can diabetes be managed effectively?
@@
Instruction:
- Answer the query using the relevant documents and the previous response. - Act like a RAG system: provide relevant, informative, and context-aware responses tailored to the user’s query intent. - Try to only answer the current utterance and explicitly do provide answers in the response, that might answer the provided upcoming utterances. - Limit your answer to no more than 150 words. - Focus on key points and avoid unnecessary repetition. - Please do answer formally and don’t use phrases like ’Okay, here’s a response to “Diabetes”’ acting as a RAG system. Answer:
Figure 2: Prompt template for generating the synthetic responses of the conversational system for Task B. Black parts show the unchanging structural and instructional parts, purple components depict the variables inserted depending on sessions and utterances.
@@ 3.2 Session Extraction from CORE Logs
Sessions were reconstructed from CORE [27] log files using a heuristic approach to group queries into sessions: starting from the last query in a potential session, queries occurring within a time window of -10 to +5 minutes were considered candidates for inclusion. Queries were added to a session if their cosine similarity, calculated using embeddings generated by the SentenceTransformer model all-MiniLM-L6-v2, with the current session queries, was greater than or equal to 0.1. All reconstructed sessions were independently manually reviewed by two people to ensure coherence and plausibility. Query sessions that contained fewer than three reformulations or did not originate from a plausible information need were excluded.
@@ 3.4 Submitted Runs
For the Micro-Shared Task we obtained 62 runs over all tasks in total and three accompanying lab notes detailing information on these runs by groups from CIR [10], Webis [21] and THM [19]. In general, we distinguish between runs that were composed (semi-) manually, through an LLM, as well as rule-based ones. We further differentiate LLM-based runs as ones composed by prompting the LLM to behave as a persona alone, through a combination of tuning and prompting and others. Table 2 provides the number of runs of different types from each group for the respective tasks.
@@ 4 Assessing Simulator Fidelity
Evaluating the quality and validity of a simulator in query prediction tasks is challenging, and the choice of suitable measures remains an open question. In this work, our methodology shifts the focus from a simulator’s effectiveness in retrieval tasks to its reproduction quality—that is, how well it replicates the authentic user behavior observed in real interaction logs. To assess this, we employ a complementary set of string-based and system-based similarity measures. Each measure is designed to capture a different aspect of simulator fidelity by quantifying the degree to which simulated queries match actual user inputs. In addition to the following measures presented in this paper, a broader overview of suitable measures was conducted in the work of Kruff et al in an additional comprehensive study [28].
@@ 4.1 Measures
Semantic Similarity. Semantic similarity assesses how close the predicted queries are in meaning to the original user queries. We computed the average cosine similarity between sentence embeddings of the original query $q_{i}^{true}$ and the $Q$ candidate queries $q_{i,j}$ over $N$ sessions. We then bounded the similarity values from [1,1][-1,1] to $[0,1]$ for better visual comparability with the other measures.
$\bar{S}=\frac{1}{|N|}\sum_{i=1}^{N}\left(\frac{1}{Q}\sum_{j=1}^{Q}\text{cosine}(q_{i}^{true},q_{i,j})\right)$
@@
Redundancy. To evaluate novelty and diversity, we introduced a Redundancy measure, which calculates the Jaccard similarity between all $Q$ candidate queries or utterances within a session averaged across all $N$ sessions:
$\bar{R}=\frac{1}{|N|}\sum_{i=1}^{N}\left(\frac{2}{Q(Q-1)}\sum_{1\leq j<q\leq Q}\text{Jaccard}(q_{i,j},q_{i,k})\right)$
Low redundancy indicates that a simulator produces multiple distinct candidates that remain semantically similar to the original query, whereas high redundancy signals minimal variation. Redundancy is not intended as a stand-alone metric; rather, it is defined for its role in the Rank-Diversity Score described below. It serves to penalize simulators that produce candidates with only minor variations, while giving simulators the opportunity to also predict queries exhibiting a degree of in-session topic drift.
SERP Overlap. To capture system-level effects, we measured the average overlap between the search engine results retrieved for the original query and the $Q$ candidates over the $N$ sessions:
$\bar{O}=\frac{1}{|N|}\sum_{i=1}^{N}\left(\frac{1}{Q}\sum_{j=1}^{Q}\text{overlap}(q_{i}^{\text{true}},q_{i,j})\right)$
Overlap denotes the fraction of shared documents in the top-10 ranked results by a fixed retrieval system. In our experiments, queries were executed using BM25. This metric assumes access to a search engine and corpus on which both original and candidate queries can be run, and is therefore dependent on the underlying retrieval model and index. The cutoff was set with reference to typical precision scores and the typical length of a webpage in the absence of pagination.
Accordingly, SERP overlap should be interpreted as a system-dependent indicator of how query reformulation affects downstream retrieval behavior, complementing string- and meaning-based similarity metrics rather than serving as a standalone effectiveness measure.
@@
$\text{RDS}_{\cos\geq 0.7}=\frac{1}{|N|}\sum_{i\in N}\left(\sum_{q\in Q_{i}}\frac{1}{\text{rank}_{q}^{\cos\geq 0.7}}\right)$
RDS is computed over all sessions $N$ and the list of candidates $Q_{i}$ generated for each session $i$. It rewards simulators that rank high-quality, diverse candidate queries or utterances at the top while penalizing poor ordering and runs submitting fewer than the required ten candidates, as this lowers their potential multiplier. Originally introduced by Carbonell and Goldstein [11], MMR provides a conceptual foundation for our measure.
@@ 5 Reflections on the Shared Task
The considerable effort invested in preparing the study environment, including comprehensive setup instructions and video guides, proved successful. None of the participating teams required substantial support from the organizers. The smooth onboarding process was a key logistical success, ensuring that the setup can be easily reproduced in future studies.
The shared task and its resulting dataset provide a valuable first step toward systematically exploring user simulation in interactive search. However, the process also highlighted several key challenges and limitations. A significant finding, particularly for Task B (conversational session simulation), was that the simulated utterances remained largely “query-like” and lacked the natural verbosity typical of this setting. This points to a major area for future improvement in simulator design. Furthermore, the semi-synthetic nature of the dataset highlighted both possibilities and constraints, offering opportunities for experimentation while revealing areas where realism could be improved. Finally, the task underscored that selecting appropriate evaluation measures is still an open and underexplored challenge for the community.